HETAV PATEL

Grade 11,

IF FORCE F=2I+3J+2K and DIsplacement r=3i+4j+5k.find the work done by force.

IF FORCE F=2I+3J+2K and DIsplacement r=3i+4j+5k.find the work done by force.

Grade:11

4 Answers

Nishant Vora IIT Patna
8 years ago
Dear student,

Work done= force dot displacement

you just need to take dot product
ADARSH REDDY
8 years ago
DEAR FRIEND,
WORK DONE =DOT PRODUCT OF FORCE AND DICPLACEMENT
SO,W=F.r
so (2i+3j+2k).(3i+4j+5k)
=6+12+10=28
so work done is 28 joules
thanks
 
SHAIK HAFEEZUL KAREEM
8 years ago
Dear student,Work done= force dot displacementyou just need to take dot product PLEASE APPROVE MY ANSWER
MAHESH
8 years ago
WE KNOW THAT WORK IS A SCALA.
THEN WORK W= F.D
GIVEN f=2I+3J+2K  ;r=3I+4J+5K
WE KNOW THAT DOT PRODUCTp { margin-bottom: 0.25cm; line-height: 120%; }

  A.B =   a1a2+b1b2+c1c2

F.D=(2×3)+(4×3)+(5×2)

    =28
THERE FORE  WORK =28J